# Decentralization of Generative AI via Mixture of Experts for Wireless Networks: A Comprehensive Survey

### v1

@@ II-D MoE with Reinforcement Learning
{tblr}
[…]
{tblr}
[…]
{tblr}
[…]
[r=1]c Vehicular Network CitySim [124] A high-precision vehicle trajectory dataset over diverse road scenarios, designed for intelligent Internet of Vehicle research. 25,000+ vehicle trajectories covering highway, urban, and intersection scenes MoE-based vehicle trajectory prediction [177]
@@ V-B4 Dataset for Wireless Networks
Datasets for wireless networks such as RadioML [122], DeepMIMO [123], 5G-NIDD [232], and CitySim [124] serve diverse wireless scenarios ranging from physical-layer signal processing to system-level intrusion detection and vehicular networking. In particular, RadioML offers labeled in-phase/quadrature (IQ) signal recordings, which are utilized by the MoE-based automatic modulation classification (AMC) model to capture distinctive signal features under varying SINR scenarios [244]. Meanwhile, DeepMIMO, built on accurate ray-tracing simulations, provides realistic mmWave and massive MIMO channels tailored to beam selection, channel estimation, and large-antenna array optimizations [245]. For network security, 5G-NIDD presents comprehensive network intrusion detection logs with fully labeled traffic traces derived from operational real-world 5G testbeds. This dataset allows MoE-based intrusion detection systems to differentiate between benign connections, newly emerging threats, and zero-day exploits, reducing false alarms and improving detection agility [111]. Additionally, the CitySim dataset focuses on connected and autonomous vehicles, delivering drone-captured vehicular trajectories that include complex mobility patterns, thereby supporting safety-oriented and MoE-based wireless research for connected and intelligent transportation systems [177].
@@ VI-A1 Lightweight and Resource-Efficient MoE Architectures
VI-A2 Dynamic Gating and Load Balancing Scheme
In dynamic wireless environments, sparse gating mechanisms are commonly employed to activate a minimal subset of experts However, sparse gating mechanisms are often susceptible to expert imbalance, where certain experts are disproportionately activated while others remain underutilized. This uneven utilization can lead to suboptimal resource efficiency and performance bottlenecks. To address this challenge, techniques such as load-balancing regularization, entropy-based sparsity constraints, and dynamic gating mechanisms (e.g., multi-objective optimization [250]) can be incorporated to promote more equitable expert allocation. These approaches strike a balance between model capability and resource utilization, enabling more effective MoE operation and sustaining real-time performance in heterogeneous wireless systems.
@@ VI-A3 Communication-Efficient MoE in Distributed Wireless Networks
In distributed wireless networks, the deployment of MoE architectures across multiple edge devices and network nodes introduces significant communication overhead and transmission delays. Efficient coordination among geographically distributed experts is critical to minimize network traffic burden, synchronize communication delays, and reduce energy consumption, especially in bandwidth-constrained wireless edge environments. Investigating efficient task partitioning, asynchronous communication protocols, and data compression and offloading techniques is essential for enabling rapid and lightweight inter-node communication. By optimizing the communication link quality and reducing redundant data flows, these approaches are capable of maintaining robust training and inference of MoE while ensuring real-time performance.

### v2

@@ V-B3 Dataset for Multimodal
    Computer Vision
    Common Image ImageNet [116] Large-scale image dataset organized by the WordNet hierarchy for visual recognition tasks 3.2 million images, 5,247 categories, 500-1,000 images per category Hard MoE [257]
    Vehicular Image nuScenes [117] Autonomous driving dataset including camera, lidar, and radar, annotated for 3D object detection and tracking 1.4 million images, 400,000 LiDAR sweeps, 1.3 million radar frames, 23 object classes LiMoE [259]
    Aerial Image DOTA [118] A large-scale aerial image dataset with oriented object annotations for real-world remote sensing applications 2806 images, 188,000 objects, 15 classes SM3Det [258]
    Multimodal
    Image-Text MS COCO [119] Natural images with human-written captions and detailed visual annotations including detection boxes, masks, and key points 330,000 images, 5 captions per image, and annotations for 80 object categories LIMOE [126]
    LAION-5B [120] A large-scale dataset includes image-text pairs in multiple languages, filtered based on CLIP similarity scores 5.85 billion image-text pairs RAPHAEL [260]
    Vedio-Text HowTo100M [121] Video–text data and ASR transcripts collected from narrated YouTube instructional videos 1.22 million videos, 136 million clip-caption pairs, 23,000+ task types Automatic speech recognition [261]
    Wireless Networks
    Physical Layer Signal Processing RadioML 2018.01a [122] Open dataset for automatic modulation classification in wireless signals 24 modulation schemes, 26 SNR levels, 1024-sample IQ data per instance MOE-AMC [178]
    Wireless Channel DeepMIMO [123] A configurable ray-tracing-based dataset for deep learning in mmWave and massive MIMO systems 12 base stations, 497,931 users, 1 blockage surface, and 2 reflector surfaces MoE-based mmWave beam selection [262]
    5G Network 5G NIDD [113] A labeled intrusion detection dataset collected from a real 5G test network, supporting ML-based security research 1.2 million flows, 9 attack types, and 1 benign traffic MoE-based intrusion detection [111]
    Vehicular Network CitySim [124] A high-precision vehicle trajectory dataset over diverse road scenarios, designed for intelligent Internet of Vehicle research. 25,000+ vehicle trajectories covering highway, urban, and intersection scenes MoE-based vehicle trajectory prediction [181]
@@ V-B4 Dataset for Wireless Networks
    Datasets for wireless networks such as RadioML [122], DeepMIMO [123], 5G-NIDD [250], and CitySim [124] serve diverse wireless scenarios ranging from physical-layer signal processing to system-level intrusion detection and vehicular networking. In particular, RadioML offers labeled in-phase/quadrature (IQ) signal recordings, which are utilized by the MoE-based automatic modulation classification model to capture distinctive signal features under varying SINR scenarios [178]. Meanwhile, DeepMIMO, built on accurate ray-tracing simulations, provides realistic mmWave and massive MIMO channels tailored to beam selection, channel estimation, and large-antenna array optimizations [262]. For network security, 5G-NIDD presents comprehensive network intrusion detection logs with fully labeled traffic traces derived from operational real-world 5G testbeds. This dataset allows MoE-based intrusion detection systems to differentiate between benign connections, newly emerging threats, and zero-day exploits, reducing false alarms and improving detection agility [111]. Additionally, the CitySim dataset focuses on connected and autonomous vehicles, delivering drone-captured vehicular trajectories that include complex mobility patterns, thereby supporting safety-oriented and MoE-based wireless research for connected and intelligent transportation systems [181].
    Datasets for wireless networks such as RadioML [122], DeepMIMO [123], 5G-NIDD [250], CitySim [124],W ViWi-Drone [264] and DroneRF [265] serve diverse wireless scenarios ranging from physical-layer signal processing to system-level intrusion detection, vehicular networking, and UAV-assisted communications. In particular, RadioML offers labeled in-phase/quadrature (IQ) signal recordings, which are utilized by MoE-based automatic modulation classification (AMC) models to capture distinctive signal features under varying SINR scenarios [178]. DeepMIMO, built on accurate ray-tracing simulations, provides realistic mmWave and massive MIMO channels tailored to beam selection, channel estimation, and large-antenna array optimizations [262]. For network security, 5G-NIDD presents comprehensive intrusion detection logs with fully labeled traffic traces derived from operational real-world 5G testbeds. This dataset allows MoE-based systems to differentiate between benign connections, newly emerging threats, and zero-day exploits, thus reducing false alarms and improving detection agility [111]. CitySim focuses on connected and autonomous vehicles, delivering drone-captured vehicular trajectories with complex mobility patterns to support safety-oriented and MoE-based research for intelligent transportation systems [181]. Additionally, ViWi-Drone and DroneRF provide realistic air-to-ground and UAV-to-UAV channel measurements, enabling MoE-based models to address dynamic beam prediction, link adaptation, and mobility-aware resource allocation in low-altitude network scenarios [266].
@@ VI-A1 Lightweight and Resource-Efficient MoE Architectures
    VI-A2 Dynamic Gating Mechanisms
    Dynamic gating represents a promising direction for enhancing the adaptability and efficiency of MoE frameworks in wireless communication systems. Future developments may focus on context-aware gating, where expert selection is conditioned on environmental variables such as user location, mobility patterns, spectrum availability, and signal-to-noise ratio, enabling adaptive prioritization of experts under congestion or energy constraints [270]. Reinforcement learning-based gating can further optimize expert activation by dynamically adjusting policies according to real-time feedback from network performance metrics such as throughput, latency, and energy consumption [271]. In addition, multi-task and multi-modal adaptive gating can facilitate shared experts across heterogeneous tasks and modalities while maintaining specialization [272]. For example, in ISAC scenarios, expert participation can be adapted depending on whether the task involves spectrum allocation, beamforming, or target detection. Finally, meta-gating and other self-evolving architectures, empowered by meta-learning, allow the gating policy itself to adapt efficiently to new operational contexts without retraining from scratch [66]. Collectively, these techniques can yield more intelligent and responsive MoE systems, particularly under the stringent performance and adaptability requirements anticipated in 6G and beyond.
@@ VI-A3 Communication-Efficient MoE in Distributed Wireless Networks
    Real-time processing is a critical requirement for distributed MoE systems deployed in time-sensitive applications such as vehicular networks and ISAC systems. However, the inherent latency of inter-agent coordination, expert selection, and output aggregation poses significant challenges to meeting stringent timing requirements. Emerging 6G technologies offer promising solutions to mitigate these constraints. Terahertz (THz) communications and extremely-large scale MIMO (XL-MIMO) provide ultra-high data rates and ultra-low transmission delays, enabling near-instantaneous exchange of gating signals and expert outputs. Cell-free massive MIMO eliminates the need for frequent handovers and supports seamless expert migration across distributed nodes, thereby reducing coordination delays in highly dynamic network topologies. Furthermore, advanced ISAC-enhancement techniques, such as joint beamforming for sensing and communication, wideband sensing, and predictive channel estimation, improve real-time responsiveness by delivering rapid environmental awareness and enabling context-driven expert activation. Leveraging these technologies allows distributed MoE frameworks to reduce expert selection latency, accelerate inter-agent communication, and deliver inference results within the strict quality requirements of 6G-enabled mission-critical services.